• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jroche

    Log entry time 15:56:55 on October 9,2009

    Entry number 294026

    keyword=day shift summary

    Day shift summary


    9:05: MCC takes the beam away for optics investigation in Hall B.
    9:25: Beam is back, we took left hrs data to check spot size (run 25000): fine, roughly 5.5*6mm.
    10:50: MCC takes the beam away for try to improve beam delivery in Hall B.

    - Lumi widths are somewhat larger than called for (know issue): the widths are now 1300 ppm when they should be 1100ppm
    - Compton rates were too high last night rendering it un-operable. One of the obejctifs for today was to allow for tuning

    List of runs (K good pairs)
    13094 -- 36
    13095 -- 23 -- 59
    13096 -- 23 -- 82
    13097 - 13103 pedestal runs (no beam)